# Are LLMs Better than Reported? Detecting Label Errors and Mitigating Their Effect on Model Performance

### v1

@@ C.1 Fine-tuning
Hardware.  For the finetuning of DeBERTa models, both the base pre-trained model, and the NLI model which is in the same size, in subsection 6.1, we used 2 Quadro RTX6000 (24GB) GPUs.
@@
Implementation.  We used HuggingFace trainer with early stopping of 4 epochs. The finetuning procedure includes splitting the training set into train and validation sets (where validation size is 25% and train 75%), fine-tuning on the train set, and choosing the best checkpoint based on the validation ROC AUC. We ran all experiments on five different seeds, affecting also the train-validation split and the random set chosen for ablation. We fine-tuned all variants with the same hyperparameters, determined by the best performing on the no-manipulation baseline. This includes 30 epochs at most, batch size of 16, learning rate of 5e-5 and weight-decay of 0.03. The rest were set as the trainer and model default.
@@
Additional Experiments.  The left plot in Figure 11 presents the same experiment discussed in subsection 6.1, but starting from the pre-trained DeBERTa-v3-base. Same trends applies here, where our LLM-confidence-based manipulations of either flipping or filtering flagged examples outperforms the baselines.
@@ C.2 Model Evaluation
Appendix D Statistical Analysis
D.1 Clopper-Pearson
As mentioned in subsection 4.1, we we employed the Clopper-Pearson exact method (Clopper & Pearson, 1934) to construct a 95% confidence interval for the binomial proportion, adjusted by a finite population correction (FPC). As we only have a subset of examples we re-annotated by LLMs or experts, we can not precisely determine what is the error rate in the full dataset, but only construct a confidence interval based on the re-annotated subset. The Clopper-Pearson method provides an exact confidence interval for a binomial proportion, which means it gives a reliable estimate even with small sample sizes. By applying the finite population correction (FPC), we adjust the interval because our sample is drawn from a limited population. This adjustment helps refine the estimate by taking into account the size of the overall dataset compared to the sample.
For your reference, the sizes of the complete datasets are provided in Table 7.
Dataset Task % pos Subset Size Full Dataset Size MNBM Summarization 10.6 640 2500 BEGIN Dialogue 38.7 640 836 VitaminC Fact Verification 52.5 640 63504 PAWS Paraphrasing 44.3 640 8000
Table 7: General datasets information.
D.2 Bootstrap sampling
In subsection 4.1, we use bootstrap sampling to provide confidence intervals for each bin. Unlike the method in subsection D.1, we do not make claims about the entire dataset, but rather focus on the re-annotated subset we possess. To achieve this, we perform 100 bootstrap samples from the empirical distribution of each bin, sampling with replacement. We then measure the agreement between the experts’ resolutions and the LLM annotations, compared to its agreement with the original label.
Appendix E Label Errors
Table 8 demonstrates one example per dataset, in which the original label is in fact an error, the LLM prediction marked it as a candidate, and the expert annotators determined the correct gold label.
@@ Appendix E Label Errors
Explanation: The generated text incorrectly states ”twin father” instead of ”twin” which is not the same, and does not even make much sense in English.
Dataset: MNBM Grounding: The John Deere tractor was pulled over by officers in the village of Ripley and had two other males on board. The vehicle had been seen in nearby Harrogate at about 05:00 GMT with no headlights on. Police said the driver had no licence, was not insured and did not have permission from the tractor’s owner. The vehicle was seized, with the three due to be interviewed by officers. Posting on Twitter, Insp Chris Galley said: ”A strange end to a night shift. 15-year-old lad driving a tractor as a taxi for his drunk mates.” Generated Text: a 15-year-old boy has been stopped by police after being seen driving a taxi on a night taxi. Original Label: 1           LLM $p$: 0.19           Gold Label: 0
Explanation: The generated text claims that the 15-year-old boy was ”driving a taxi on a night taxi”, contradicting the grounding in which it was claimed that the boy was driving a tractor as a taxi

### v2

@@ F.1 Fine-tuning
For the finetuning of DeBERTa models, both the base pre-trained model, and the NLI model which is in the same size, in subsection 7.1, we used 2 Quadro RTX6000 (24GB) GPUs.
Implementation.
@@
We finetuned starting from two base models: DeBERTa-v3 444microsoft/deberta-v3-base, and a fine-tuned version of it on classic NLI datasets  [^5] MoritzLaurer/DeBERTa-v3-base-mnli-fever-anli. We used HuggingFace trainer with early stopping of 4 epochs. The finetuning procedure includes splitting the training set into train and validation sets (where validation size is 25% and train 75%), fine-tuning on the train set, and choosing the best checkpoint based on the validation ROC AUC. We ran all experiments on five different seeds, affecting also the train-validation split and the random set chosen for ablation. We fine-tuned all variants with the same hyperparameters, determined by the best performing on the no-manipulation baseline. This includes 30 epochs at most, batch size of 16, learning rate of 5e-5 and weight-decay of 0.03. The rest were set as the trainer and model default.
Additional Experiments.
@@
The left plot in Figure 14 presents the same experiment discussed in subsection 7.1, but starting from the pre-trained DeBERTa-v3-base. Same trends applies here, where our LLM-confidence-based manipulations of either flipping or filtering flagged examples outperforms the baselines.
@@ F.2 Model Evaluation
Fine-Tuning vs. Zero-Shot
Interestingly, the overall trend of improved performance on the corrected labels does not hold for the DeBERTa-based fine-tuned models. Unlike the LLMs, which are prompted in a zero-shot setting, the fine-tuned models are trained on the original dataset, which contains label errors. As a result, the LLMs demonstrate better generalization, while the fine-tuned models may overfit to the noise in the training data. A plausible explanation for this reversed trend lies in the distributional prior learned from the training set. In the original dataset, labels of 0 (inconsistent) are more frequent than in the corrected gold set. For example, among examples where the original and gold labels agree, the proportion of 1 (consistent) labels is 36%, and the model (DeBERTa-v3-base predicts 1 in 35% of those cases. In contrast, among examples where the labels disagree, the gold rate of 1 is 58%, yet the model predicts 1 in only 36% of the cases. This pattern suggests that the model has learned a skewed prior from the flawed dataset, underestimating the likelihood of the consistent class, particularly in cases that were originally mislabeled. Similar percentages are observed for the NLI model as well.
Appendix G Statistical Analysis
G.1 Clopper-Pearson
As mentioned in subsection 5.1, we employed the Clopper-Pearson exact method (Clopper and Pearson, 1934) to construct a 95% confidence interval for the binomial proportion, adjusted by a finite population correction (FPC). As we only have a subset of examples we re-annotated by LLMs or experts, we can not precisely determine what is the error rate in the full dataset, but only construct a confidence interval based on the re-annotated subset. The Clopper-Pearson method provides an exact confidence interval for a binomial proportion, which means it gives a reliable estimate even with small sample sizes. By applying FPC, we adjust the interval because our sample is drawn from a limited population. This adjustment helps refine the estimate by taking into account the size of the overall dataset compared to the sample.
G.2 Bootstrap sampling
In subsection 5.1, we use bootstrap sampling to provide confidence intervals for each bin. While not necessarily the first to introduce it, Xia et al. (2012) explored bootstrap confidence intervals on ROC AUC. Unlike the method in Appendix G.1, we do not make claims about the entire dataset, but rather focus on the re-annotated subset we possess. To achieve this, we perform 100 bootstrap samples from the empirical distribution of each bin, sampling with replacement. We then measure the agreement between the experts’ resolutions and the LLM annotations, compared to its agreement with the original label.
Appendix H Label Errors
Table 6 demonstrates one example per dataset, in which the original label is, in fact, an error, the LLM prediction marked it as a candidate, and the expert annotators determined the correct gold label.
@@ Appendix H Label Errors
Explanation: The generated text incorrectly states "twin father" instead of "twin" which is not the same, and does not even make much sense in English.
Dataset: MNBM Grounding: The John Deere tractor was pulled over by officers in the village of Ripley and had two other males on board. The vehicle had been seen in nearby Harrogate at about 05:00 GMT with no headlights on. Police said the driver had no licence, was not insured and did not have permission from the tractor’s owner. The vehicle was seized, with the three due to be interviewed by officers. Posting on Twitter, Insp Chris Galley said: "A strange end to a night shift. 15-year-old lad driving a tractor as a taxi for his drunk mates." Generated Text: a 15-year-old boy has been stopped by police after being seen driving a taxi on a night taxi. Original Label: 1      LLM $p$: 0.19      Gold Label: 0
Explanation: The generated text claims that the 15-year-old boy was "driving a taxi on a night taxi", contradicting the grounding in which it was claimed that the boy was driving a tractor as a taxi